Please help me with these questions.

1). Let f(x) = 3x2 + 2 and g(x) = 3x − 3. Find the function. (g ∘ f)(x)

2). Write an equation in standard form of the circle described.

Ends of diameter at (−2, −2) and (4, 6)
Answer #1

Hope this helps you! Thank you. criven that kn) = 32+2 gen) = 3x-3. Now, (gof) m) = g( Fon)) 8 (3x+2) = 3 (3 042-3 =gart. 6-3If you don't understand please comment